Larry is now twice as old as Bill.Five years ago Larry was athree times as old as Bill.Find their present ages.
Answer by jorel1380(3719)

Let Larry equal l,Bill equal b:
l=2b
l-5=3(b-5)
2b-5=3b-15
10=b
20=l